### v1

# Seo Hyun Kim1     Kai Tzu-iunn Ong1 Taeyoon Kwon1     Namyoung Kim1     Keummin Ka1     SeongHyeon Bae1 Yohan Jo2     Seung-won Hwang2     Dongha Lee1     Jinyoung Yeo1

Abstract
Large language models (LLMs) are capable of processing lengthy dialogue histories during prolonged interaction with users without additional memory modules; however, their responses tend to overlook or incorrectly recall information from the past. In this paper, we revisit memory-augmented response generation in the era of LLMs. While prior work focuses on getting rid of outdated memories, we argue that such memories can provide contextual cues that help dialogue systems understand the development of past events and, therefore, benefit response generation. We present [Uncaptioned image]Theanine, a framework that augments LLMs’ response generation with memory timelines – series of memories that demonstrate the development and causality of relevant past events. Along with Theanine, we introduce TeaFarm, a counterfactual-driven question-answering pipeline addressing the limitation of G-Eval in long-term conversations. Supplementary videos of our methods and the [Uncaptioned image]TeaBag dataset for TeaFarm evaluation are in https://theanine-693b0.web.app/.
[Uncaptioned image]Theanine: Revisiting Memory Management in Long-term Conversations with Timeline-augmented Response Generation
Seo Hyun Kim1thanks: These two authors are the co-first authors and both have the right to list their names first on their CVs.     Kai Tzu-iunn Ong1 Taeyoon Kwon1thanks: T Kwon, N Kim, and K Ka are the co-second authors.     Namyoung Kim1     Keummin Ka1     SeongHyeon Bae1 Yohan Jo2     Seung-won Hwang2     Dongha Lee1     Jinyoung Yeo1 1Yonsei University, 2Seoul National University
1 Introduction
Recently, the impressive performance of large language models (LLMs) has caused a trend of having longer conversations with them across multiple sessions (Fui-Hoon Nah et al., 2023). Such prolonged interactions require the models to memorize previous events or speaker information, tailoring the responses accordingly to keep up user interest (Adiwardana et al., 2020). While the large context windows of LLMs [^1] Recent LLMs, such as GPT-4 (Achiam et al., 2023) and LLama3 (MetaAI, 2024), have contexts windows of 128,000 and 1,040,000 tokens, respectively. allow them to process the entire dialogue history for response generation (RG), this often leads to biased attention toward the latest user utterance, ignoring relevant contexts from the past (Liu et al., 2024), as presented in Figure 1 (a).
Figure 1: Empirical examples of failed responses due to (a) bias to the latest input (b) the absence of an important past event (“afraid of cruise ships”) on the timeline. (c) shows a response augmented with the memory timeline.
Figure 2: The overview of Refer to captionTheanine. Left: Linking new memories to the memory graph after finishing a dialogue session; Right: Memory timeline retrieval, refinement, and response generation in a new dialogue session.
One straightforward solution is to condense past conversations into summarized memories and retrieve them to augment response generation in later sessions (Xu et al., 2022a, b; Lu et al., 2023), but the growing span of memories can hinder the quality of memory retrieval as dialogue sessions accumulate. Although this can be mitigated by updating the older memories (Bae et al., 2022; Zhong et al., 2024; Kim et al., 2024a), we observe severe information loss. As illustrated in Figure 1 (b), an earlier memory on the timeline, which contains an important speaker persona (i.e., “afraid of ships”), is removed during memory update (thus can not be retrieved), yielding a suboptimal response.
Motivated by these, in this paper, we revisit memory-augmented response generation for long-term conversations and address the aforementioned bottlenecks. Specifically, we focus on using the timeline built with relevant memories to augment response generation. To this end, we present [Uncaptioned image]Theanine, [^2] L-theanine is an amino acid found in green tea that has been linked to memory improvement (Nguyen et al., 2019). a framework of timeline-augmented chain-of-thought reasoning for response generation in long-term conversations.
Inspired by how humans naturally link new memories to existing memories of relevant events and retrieve them as a whole (Bartlett, 1995), Theanine manages memories with a graph structure, i.e., the memory graph, where memories are linked based on how they relate to each other (Phase I). Intuitively, this can be done by linking each memory to others that have high text similarity. Yet, we find this naive approach not much beneficial to response generation, due to the plain connection between memories (e.g., “this happened $\rightarrow$ that occurred”). To resolve this, we leverage an LLM to dynamically link memories based on their temporal and cause-effect commonsense relations (Hwang et al., 2021) via our designed algorithm. Next, in memory retrieval for generating a new response (Phase II-1), instead of solely retrieving top-$k$ memories using current conversation as query, we retrieve the whole memory timeline representing the development of relevant events. Then, to tackle the discrepancy between off-line memory construction and online deployment (Phase II-2), Theanine refines the retrieved timeline such that it provides tailored information for the current conversation. Finally, in response generation (Phase III), we leverage LLMs’ chain-of-thought (CoT) reasoning ability (Wei et al., 2022) to reason over current conversation and refined timelines, conclude useful information, and generate the final response. Our contributions are three-fold:
  • We revisit memory-augmented response generation in long-term conversations with a novel timeline-augmented framework, [Uncaptioned image]Theanine. In both LLM-based (i.e., G-Eval) and human evaluation, Theanine demonstrates (i) better responses that are detailed/non-generic and (ii) efficiency in properly referencing past conversations.
  • The lack of ground-truth mapping between ongoing conversations (queries) and memories for retrieval poses a challenge in evaluating memory-augmented dialogue systems. We present TeaFarm, a counterfactual-driven question-answering pipeline, that assesses systems in referencing past conversations, which does not require human intervention.
  • We further curate [Uncaptioned image]TeaBag, the counterfactural question-answer data for TeaFarm evaluation on two widely used benchmark datasets of long-term conversations: Multi-Session Chat (MSC) (Xu et al., 2022a) and Conversation Chronicles (CC) (Jang et al., 2023).
2 Methodologies
We present [Uncaptioned image]Theanine (Figure 2), a framework that resembles the mechanism of how humans store and retrieve memories for conversations:
2.1 Memory Graph Construction (Phase I)
To replicate how humans construct memories, we approach memory management based on a graph structure, i.e., the memory graph $G$:
$G=(V,E)$ (1)
$V=\{m_{1},m_{2},...,m_{|V|}\}$ (2)
$m=(\textit{event},\textit{time})$ (3)
$E=\{\langle m_{i},r_{ij},m_{j}\rangle|m_{i},m_{j}\in V\land r_{ij}\in R\}$ (4)
$R=\{\texttt{Cause},\texttt{Reason},\texttt{Want},...,\texttt{SameTopic}\}$ (5)
In $G$, vertices $V$ are memories $m$ summarized from the conversations. Each memory $m=(\textit{event},\textit{time})$ consists of an event [^3] In this work, “event” denotes information perceived by the dialogue system, including things done/said by speakers and the acknowledgement of speaker personas. and the time it is formed (summarized). Each directed edge $e\in E$ between two connected $m$ indicates their temporal order and their cause-effect commonsense relation $r\in R$:
At the end of dialogue session $t$, [Uncaptioned image]Theanine starts linking each new memory $m_{new}$ summarized from session $t$ to the memory graph $G^{t}$.
Phase I-1: Identifying associative memories for memory linking.
Following how humans link new memories to existing ones that are related to a similar event/topic, i.e., the associative memories, Theanine starts by identifying these associative memories from the memory graph $G^{t}$.
Formally, given a newly-formed memory $m_{new}$ waiting to be stored, the associative memories $M_{a}$ of $m_{new}$ is defined as the set of $m_{i}\in G^{t}$ having top-$j$ text similarity with $m_{new}$ (i.e., $|M_{a}|=j$).
Phase I-2: Relation-aware memory linking.
Intuitively, we can link $m_{new}$ to $m\in M_{a}$ using edges that indicate their text similarity and chronological order, we find such simplified connection (e.g., “this happened $\rightarrow$ that similar event occurred”) can yield a context-poor graph that does not help response generation much (Section 4).
Humans, on the other hand, interpret events by considering the relation between them, such as “how does an event affect the other?” or “why did this person make that change?”.
Therefore, we adopt a relation-aware memory linking, where each edge between two memories is encoded with their cause-effect commonsense relation $r\in R$, along with their temporal order. These relations are adopted from Hwang et al. (2021) and modified by us, including HinderedBy, Cause, Want, and 4 more (Appendix A.1).
We start by determining the relation between $m_{new}$ and each associative memory. Formally, for each pair of $m_{new}$ and $m\in M_{a}$, the LLM assigns a relation $r\in R$ based on their event, time and their origin conversations:
$M_{a}^{*}=\{m_{i}\in M_{a}\mid\Upsilon(m_{i},m_{new})\in R\}$ (6)
where $\Upsilon(\cdot,m_{new})\in R$ indicates that the given memory is assigned with an $r\in R$ with $m_{new}$, [^4] Limited by the performance of retrievers, it is possible that an $m\in M_{a}$ does not have a relation, other than just textual overlap, with $m_{new}$. We address this by allowing the LLM to output “None”. and such assigned memories are defined as $M_{a}^{*}$.
Figure 3: Locating memories to be linked to $m_{new}$.
We then proceed to link $m_{new}$ to the graph, a straightforward approach is linking $m_{new}$ to all $m\in M_{a}^{*}$. Yet, we find it algorithmically inefficient (Figure 7). Thus, we selectively link $m_{new}$ to only those $m\in M_{a}^{*}$ that are recent. To identify those recent memories, we first locate every connected component $C_{i}\subset G^{t}$ that contains at least one $m\in M_{a}^{*}$, as shown in Figure 3 (a) and (b):
=$\mathbb{C}=\{C_{i}\subset G^{t}\mid\mathtt{V}(C_{i})\cap M_{a}^{*}\neq\emptyset\ \}$ (7)
where $\mathbb{C}$ is the collection of those $C$ and $\mathtt{V}(\cdot)$ represents “vertices in”. Now, we link $m_{new}$ to the most recent $m\in M_{a}^{*}$ in each $C_{i}\subset\mathbb{C}$ (Figure 3 (c)). Therefore, the memories $M_{linked}$ that are linked to $m_{new}$ is defined as follows:
$M_{linked}=\{\Omega(\mathtt{V}(C_{i})\cap M_{a}^{*})\mid C_{i}\subset\mathbb{C}\}$ (8)
where $\Omega(\cdot)$ indicates “the most recent memory in”. Linking all memories from session $t$ to $G^{t}$, we then obtain a new memory graph $G^{t+1}$. The pseudo algorithm for Phase I is in Algorithm 1.
2.2 Timeline Retrieval and Timeline Refinement (Phase II)
Drawing motivation from Figure 1, [Uncaptioned image]Theanine augments response generation (RG) with timelines of relevant events to address the information loss caused by conventional memory management (Xu et al., 2022a; Bae et al., 2022). With $G^{t+1}$, Theanine performs these steps for RG in session $t+1$:
Preparation: Top-k memory retrieval.
During the conversation, using the current dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ of $n$ utterances $u$ as query, we first retrieve top-$k$ memories $M_{re}=\{m_{re1},...,m_{rek}\}$.
Phase II-1: Retrieving raw memory timelines.
After that, we wish to also access memories that are centered around relevant events. Formally, given an $m_{re}\in M_{re}$, we further retrieve the connected component $C_{re}\subset G^{t+1}$ that contains $m_{re}$.
Since this collection of memories (i.e., $C_{re}$) can be “tangled up” together (i.e., connected in a complex manner) due to the graph structure, we proceed to untangle it into several memory timelines, each representing a series of events about $m_{re}$ that starts out similarly yet branches into slightly different development. For that, we first locate the earliest memory in $C_{re}$ as a starting point $m_{start}$ for all timelines, as shown in Figure 4 (left).
$m_{start}=\Theta(\mathtt{V}(C_{re}))$ (9)
where $\Theta$ indicates “the oldest memory in”
Figure 4: Extracting raw memory timelines $\tau$ from the retrieved connected component $C_{re}$.
Next, starting from $m_{start}$, we untangle the memories by tracing through the future direction and extract every possible linear graph containing $m_{re}$ (two in Figure 4) from $C_{re}$, until reaching an endpoint $\tau[-1]$ with an out-degree of 0 (i.e., $deg^{+}(\tau[-1])=0$, which means no directed edge goes out from it). Each of them is considered a raw memory timeline $\tau$, demonstrating a version of the evolution of $m_{re}$ and its relevant events:
$\begin{split}\mathcal{T}={}&\{\tau\subset C_{re}\mid\tau\smash{{\text{ is a directed linear}}}\\ &\smash{{\text{ graph s.t. }}}m_{start},m_{re}\in\tau\ \\ &\land deg^{+}(\tau[-1])=0\}\end{split}$ (10)
We then sample $n$ raw timelines $\tau$ from $\mathcal{T}$ [^5] Based on our experiments, we empirically set $n$ to 1. Repeating Phase II-1 for all retrieved top-$k$ memories, we collect a set of retrieved raw memory timelines $\mathbb{T}=\cup\,\mathcal{T}$, where $|\mathbb{T}|=k^{*}n$.
Phase II-2: Context-aware timeline refinement.
Although we have constructed the memory graph using temporal and commonsense relations to improve its informativeness, directly applying the retrieved timelines for response generation can be suboptimal (RQ3, Section 4), because graph construction does not take the current conversation into consideration, i.e., they are constructed off-line.
In this phase, [Uncaptioned image]Theanine tackles such a discrepancy between off-line memory construction and online deployment (i.e., an ongoing current conversation) via a context-aware timeline refinement. Motivated how LLMs can self-refine their previous generation (Madaan et al., 2024). We leverage LLMs to refine raw timelines into a rich resource of information crafted for the current conversation, by removing redundant information or highlighting information that can come in handy. Formally, given the current dialogue $\mathcal{D}$ and the retrieved raw memory timelines $\mathbb{T}$, the LLM tailors all $\tau\in\mathbb{T}$ into refined timelines $\mathbb{T}_{\Phi}$:
$\mathbb{T}_{\Phi}=\{\underset{\tau_{\Phi}}{\text{argmax}}\,P_{\text{LLM}}(\tau_{\Phi}|\mathcal{D},\tau)\mid\tau\in\mathbb{T}\}$ (11)
All refined timelines $\mathbb{T}_{\Phi}$ are then used to augment the response generation. We provide the pseudo algorithm for Phase II in Algorithm 2.
2.3 Timeline-augmented Response Generation (Phase III)
Finally, we move on to response generation. Here, [Uncaptioned image]Theanine leverages LLMs’ chain-of-thought (CoT) reasoning to discover user intention from the current conversation (Chae et al., 2023) and selectively extract useful information provided in the refined timelines. Formally, given the current dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ and refined timelines $\mathbb{T}_{\Phi}$, the LLM sequentially generates a CoT rationale $\mathcal{R}$ demonstrating how it comprehends $\mathcal{D}$ and $\mathbb{T}_{\Phi}$, and the next response $\bar{u}_{t+1}$:
¯$\bar{\mathcal{R}}=\underset{\mathcal{R}}{\text{argmax}}\,P_{\text{LLM}}(\mathcal{R}|\mathcal{D},\mathbb{T}_{\Phi})$ (12)
$\Rightarrow\bar{u}_{n+1}=\underset{u_{n+1}}{\text{argmax}}\,P_{\text{LLM}}({u}_{n+1}|\mathcal{D},\mathbb{T}_{\Phi},\bar{\mathcal{R}})$ (13)
where $\Rightarrow$ denotes sequential text generation.
3 Experimental Setups
3.1 Datasets of Long-term Conversations
There are few datasets of long-term conversations. First, Multi-Session Chat (MSC) (Xu et al., 2022a), which is built upon the conversations from Persona-Chat (Zhang et al., 2018) by extending their follow-up conversations throughout multiple sessions. Soon after MSC, DuLeMon (Xu et al., 2022b) and CareCall (Bae et al., 2022) are proposed for long-term conversations in Mandarin and Korean, respectively. Recently, Jang et al. (2023) release a new multi-session datasets, Conversation Chronicles (CC). Different from MSC, speakers in CC are augmented with diverse relationships, such as “employee and boss”. Apart from the above open-domain datasets, there is also Psychological QA, [^6] https://www.xinli001.com/ which addresses long-term conversation in Mandarin regarding clinical scenarios.
In this work, we adopt MSC and CC for evaluation. Since we focus on English conversations, we do not include DuleMon, CareCall, and Psychological QA. We leave such multilingual settings and clinical scenarios to future work.
Datasets: Multi-session Chat (MSC) & Conversation Chronicles (CC)
Dialogue Session: Session 3 Session 4 Session 5 Avg.
[Uncaptioned image]Theanine Theanine’s Winning Percentage
          vs. Spe Mem Cst Spe Mem Cst Spe Mem Cst Spe Mem Cst
All Dialogue History 71% 82% 77% 73% 81% 77% 72% 79% 77% 72% 81% 77%
All Memories and $\mathcal{D}$ 69% 64% 66% 69% 69% 68% 70% 68% 68% 69% 67% 67%
     + Memory Update 56% 57% 60% 56% 56% 58% 56% 58% 60% 56% 57% 60%
Memory Retrieval 67% 62% 68% 72% 67% 68% 72% 65% 68% 70% 65% 68%
     + Memory Update 57% 56% 59% 58% 57% 61% 55% 58% 59% 57% 57% 60%
RSum-LLM 38% 61% 69% 35% 57% 67% 37% 57% 65% 37% 58% 67%
MemoChat 60% 74% 71% 60% 69% 72% 60% 72% 71% 60% 72% 71%
COMEDY 66% 54% 78% 64% 54% 77% 62% 53% 76% 64% 53% 77%
Golden Responses 63% 76% 61% 63% 73% 61% 62% 72% 61% 63% 73% 61%
Table 1: Win rates of Theanine in head-to-head G-Eval (of all generated responses) with baselines. Spe = specificity; Mem = memorability; Cst = consistency. Color green denotes metrics regarding the reference of memories. Color gray shows where ours loses more. Underline: p-value $<0.001$. *: $0.001<$ p-value $<0.05$.
3.2 Baselines
To evaluate Theanine in long-term conversation, besides naive settings using all past dialogues or memories, we include the following baselines.
Memory retrieval. Following Xu et al. (2022a) and Xu et al. (2022b), this setting uses a retriever to retrieve memories with the current dialogue context to augment response generation.
+ Memory update. We use LLMs to perform the memory updating mechanism developed by Bae et al. (2022) to update the memory pool at the end of each session to enhance the above baselines.
RSum-LLM. Proposed by Wang et al. (2023), it uses an LLM to (i) recursively summarize and update the memory pool at the end of each session, (ii) select top memories based on the current dialogue, and (iii) generate responses.
MemoChat. Proposed by Lu et al. (2023), it uses LLMs’ CoT reasoning ability to (i) conclude important memories from past conversations in a structured topic-summary-dialogue manner, (ii) select memories, and (ii) generate responses.
COMEDY. Proposed by Chen et al. (2024b), rather than memory retrieval, it first summarizes session-level past summaries and then compresses all of them into short sentences of events, user portraits (behavior patterns, emotion, etc.) and user-bot relation, using them to augment response generation.
3.3 Models and Implementation Details
Large language models. In all experiments including baselines, we adopt ChatGPT (gpt-3.5-turbo-0125) from OpenAI (2023) for (i) memory summarization, (ii) memory update, and (iii) response generation. We set the temperature to $0.75$.
Retrievers. We use text-embedding-3-small (OpenAI, 2024b) from OpenAI to calculate the similarity of texts. For the identification of top-$j$ associative memories (Phase I-1) and top-$k$ memory retrieval (Phase II Preparation), we set $j$ and $k$ to 3. For the “Memory Retrieval” baseline, we retrieve 6 memories following Xu et al. (2022a).
Dialogue sessions. We perform evaluation using sessions 3-5 from both datasets, as all settings are almost identical before the end of session 2 (there is no memory to update before then).
4 Evaluation Scheme 1: Machine and Human Evaluation
To evaluate [Uncaptioned image]Theanine, we first start with G-Eval (Liu et al., 2023), an LLM-based framework for assessing LMs’ responses, and human evaluation. [^7] Details for G-Eval and human evaluation are in Appendix D and F. Dataset-specific results are in Appendix E. Here, we focus on the following research questions (RQs):
RQ1. Can Theanine generate context-rich responses, rather than those generic and plain?
RQ2. Can Theanine reference past conversations better than other methods?
RQ3. Do phases in Theanine make sense?
RQ4. Do humans agree with Theanine’s procedures, and how do they think of the final responses?
[Uncaptioned image]Theanine yields responses that are more detail-rich (RQ1).
In Table 1, we first assess whether generated responses are detail-rich, rather than generic and plain, i.e., specificity (Spe). In both datasets, Theanine yields significantly high win rates in the head-to-head comparison with all baselines except for RSum-LLM. We presume it is due to the fact that RSum-LLM tends to generate much longer responses (avg. 77 words per response) than other settings (ours = 32 words; avg of other baselines = 30 words).
[Uncaptioned image]Theanine utilizes past memories efficiently and correctly (RQ2).
In long-term conversations, it is crucial and necessary to further assess (i) whether the dialogue system utilizes information from the past when necessary (i.e., memorability) and make sure (ii) the response does not contradict previous conversations (i.e., consistency). We present the results in Table 1. Theanine demonstrates high win rates in the comparisons with baselines. However, COMEDY performs almost as well as us regarding memorability, which we assume is because it always applies two additional LLMs to extract and compress past conversations into concise memos before each response generation (no retrieval). However, as we expected, always referencing concise and recent user information can lead to contradiction to older information, thus clearly beaten by Theanine in terms of consistency (Cst).
Ablation results: Phases in [Uncaptioned image]Theanine matter (RQ3).
Table 2 presents the results where key phases in Theanine are removed. We observe that the removal of relation-aware memory linking (Phase I-2) and context-aware timeline refinement (Phase II-2) leads to large performance drops regarding the reference of past information (i.e., the original Theanine shows higher win rates), especially when timeline refinement is absent.
This justifies our methodological designs: (i) In graph construction, linking memories with cause-effect relation benefits RG more than naive connection using only text similarity and chronological order; (ii) By further refining memories built off-line (i.e., without considering current conversations) before RG, we can elicit contextual cues that are tailored for the current conversation.
Multi-session Chat (Session 3-5)
[Uncaptioned image]Theanine Ours’ Win Rate
          vs. Mem Cst Avg.
w/o Relation-aware Linking 62% 65% 64%
w/o Timeline Refinement 62% 69% 65%
Conversation Chronicles (Session 3-5)
w/o Relation-aware Linking 60% 65% 63%
w/o Timeline Refinement 64% 68% 66%
Table 2: Ablation results (p-values $<0.001$).
Figure 5: Results of human evaluation.
Humans think [Uncaptioned image]Theanine makes sense and properly references past conversations.
As reported in Figure 5, human judges largely agree (92%) that Theanine’s relation-aware memory linking is performed appropriately, and they all agree (100%) that timeline refinement can elicit more helpful/tailored information for response generation. Most importantly, Theanine produces responses that either entail (i.e., faithfully reflect) past conversations (34%) or, at worse, neutral statements (57%) in 91% of the cases from human standpoints, manifesting its’ efficacy in referencing past information in long-term conversations.
Examples of [Uncaptioned image]Theanine in terms of each phase are in Appendix G.
5 Evaluation Scheme 2: TeaFarm – a Counterfactual-driven Evaluation Pipeline for Long-term Conversations
Despite the success of G-Eval (Liu et al., 2023) in evaluating responses of LMs, accessing dialogue systems’ ability to utilize past memories in long-term conversation still remains challenging. This is due to the unavailability of ground-truth mapping between current conversations and the correct memories for retrieval. Although we can resolve this by feeding the evaluator LLM (e.g., GPT-4) the entire memory pool or past history and prompt it to determine whether a response correctly recalls past conversations, the evaluation can be limited by the performance of the evaluator LLM itself.
To overcome this, along with Theanine, we present TeaFarm, a counterfactual-driven pipeline for evaluating memory-augmented response generation in long-term conversations.
5.1 Testing Dialogue Systems’ Memory via Counterfactual Questions
In TeaFarm, we proceed to “trick” the dialogue systems into generating incorrect responses, and the system must correctly reference past conversations to avoid being misled by us. Specifically, we talk to the dialogue system while acting as if a non-factual statement is true (thus counterfactual). Figure 6 presents some examples of counterfactual questions and the corresponding facts.
Figure 6: Examples of counterfactual questions.
In practice, when we want to evaluate a system that has been interacting with the user for multiple sessions, we first (1) collect all past conversations and summarize them session by session. Then, we (2) feed the LLM [^8] We apply GPT-4 (gpt-4) with a temperature of $0.75$. the collected summaries in chronological order such that it can capture the current stage of each discussed event, e.g., “Speaker B has never been to Japan at this moment”, and (3) generate counterfactual questions from the perspective of both speakers (and the correct answers). After that, we (4) kick off a new conversation, chat for a while, then (5) naturally ask the counterfactual question, and (6) assess the correctness of its response. Overview illustration and prompts for TeaFarm are in Appendix B and H.
5.2 TeaFarm Results
Settings / Datasets MSC CC Avg.
Memory Retrieval 0.16 0.19 0.18
+ Memory Update 0.16 0.19 0.18
RSum-LLM 0.04 0.08 0.06
MemoChat 0.09 0.15 0.12
COMEDY 0.06 0.18 0.12
[Uncaptioned image]Theanine 0.17 0.24 0.21
      w/o Relation-aware Linking 0.06 0.13 0.10
      w/o Timeline Refinement 0.15 0.16 0.16
Table 3: TeaFarm results regarding correct/total percentages of model responses. Tested in conversations from MSC and CC with 200 counterfactual questions.
In Table 3, we present model performance in TeaFarm evaluations. Theanine outperforms all baselines, especially in CC. Ablations perform worse than original Theanine, again proving the efficacy and necessity of relation-aware linking and timeline refinement. Surprisingly, all settings result in low correct rates, qualifying TeaFarm as a proper challenge for stress-testing dialogue systems in long-term conversations.
Interestingly, baselines that utilize retrievers (just like Theanine) show superior performance than those relying on LLMs for memory selection (i.e., RSum-LLM, MemoChat, and COMEDY). [^9] Memory update does not affect Memory Retrieval’s performance. We believe it is because the counterfactual questions are made to counter the newest stage of each event. The removal of older memories thus does not have much impact. This pattern, unexpectedly, supports our idea of revisiting memory retrieval in the era of LLMs.
5.3 Dataset for TeaFarm: [Uncaptioned image]TeaBag
As a byproduct of TeaFarm, we curate [Uncaptioned image]TeaBag, a dataset for TeaFarm evaluation on MSC and CC. TeaBag consists of (1) 100 episodes of original conversations from MSC and CC (session 1-5), (2) follow-up conversations (session 6), where the speaker will naturally converse toward the counterfactual questions, and (3) the corresponding answers. TeaBag serve as a useful tool for stress-testing if a system can correctly reference past conversations. Dataset details are in Appendix C.
6 Further Discussions
Cost efficiency.
In Phase I-2, we choose to link $m_{new}$ to only the most recent associative memories in each $C_{i}$, rather than all associative memories in $M_{a}$. Figure 7 compares Theanine with Theanine-All, where $m_{new}$ is linked to all $m\in M_{a}$. We find that although Theanine-All costs 25% more, it is beaten by original Theanine. We believe this is because linking $m_{new}$ to all $M_{a}$ leads to more short timelines (e.g., an $m_{new}$ from session 4 is directly linked to an old memory from session 1), which provide less information than timelines built with relevant memories from more sessions. Regardless, it still yields lower loss rates than strong baselines. An open question here is whether we can reach a balance between cost and performance. We presume we can further improve the cost efficiency of Theanine in the long-term by training a memory linker using the collected memory-relation pairs from LLMs via knowledge distillation. We leave it to future work.
Growing span of memories.
Another inquiry left might be whether the growing span of memory would hinder memory retrieval in Theanine if there ever were hundreds of sessions. Although this can be a serious issue for conventional memory retrieval, we argue that it will be much mitigated in Theanine because: (i) We retrieve relevant memories as a whole in the form of timelines. This works as a safety net in scenarios where an important memory is missed out in top-$k$ retrieval – it might still be retrieved along with the timeline; (ii) We perform timeline refinements. Similarly, this also acts as a second insurance against mis-retrieved memories that are actually not helpful.
Figure 7: API cost and G-Eval results. Red shows baselines that lose even more than Theanine-All in the head-to-head comparisons with original Refer to captionTheanine.
7 Related Work
Long-term Conversations.
Since the release of MSC (Xu et al., 2022a), there have been many studies on long-term conversations. For instance, Bae et al. (2022) train a classifier as the memory updater to handle outdated memories in phone call scenarios. Recently, based on LLMs’ outstanding in-context learning ability, Wang et al. (2023) leverage LLMs to write, update, and select memories for response generation. Apart from LLMs’ power, human behaviors have also fostered methods for long-term conversations. For example, Zhong et al. (2024) apply humans’ forgetting curve to make memories that have been discussed stay in the memory pool for a longer time. Also drawing inspiration from how humans manage memories, our [Uncaptioned image]Theanine, on the other hand, addresses the connection between relevant events in both memory construction and retrieval, improving response quality in long-term conversations, which has never been explored to the best of our knowledge.
Memory-augmentation for personalized dialogue systems.
The trend of long-term interaction with dialogue systems promotes the adaptation of them for personalized needs (Chen et al., 2024a, c). As a pioneer, Xu et al. (2022b) train a persona extractor to create user-based memories in long-term conversations. However, training a personalized system for long-term use can be non-trivial due to the lack of data (Tseng et al., 2024). As a solution, Kim et al. (2024a) apply commonsense models and LLMs to augment existing long-term conversation datasets with high-quality persona sentences; Chen et al. (2024b) present a training-free LLM-based framework that extracts user behaviors from past conversations for personalized responses. Upon the success of LLMs, [Uncaptioned image]Theanine leverages them to build memory timelines, which represent the development of user personas/behaviors throughout the interaction, providing rich cues for personalized responses.
8 Conclusions
This work revisits memory management in multi-session/long-term conversations in the era of LLMs. Following how humans store and retrieve relevant memories as a whole, we present [Uncaptioned image]Theanine, a framework of timeline-augmented response generation, which elicits responses that are detail-rich and, more importantly, able to correctly reference previous conversations when necessary. We also propose a novel evaluation pipeline (TeaFarm) that addresses the limitation of G-Eval in long-term conversations as well as the dataset for it ([Uncaptioned image]TeaBag). We expect this paper to serve as a groundwork and inspiration for exploring more paradigms of memory construction/retrieval in dialogue systems.
9 Limitations
This work has the following limitations: First, the maximum amount of dialogue sessions in this study is limited to five due to the lack of longer open-domain English datasets of multi-session conversations. Hence, Theanine’s performance in longer conversations has not been investigated. As we mentioned in Section 6, we presume that its’ effectiveness can still hold true to some degree even in such scenarios. Yet, we do acknowledge the need to apply additional modules that directly address the growing span of dialogue history/memories. One possible way is to introduce the summarize-then-compress paradigm in COMEDY (Chen et al., 2024b) to Theanine, which further compresses session-level summaries into a combined short user/event description.
Second, retrieving and refining timelines for response generation can be less computationally efficient than conventional memory retrieval, as there might exist scenarios where timelines are not necessary, e.g., when there is no need to reference past conversations. We believe this can be mitigated by incorporating an additional selection logic/module to choose between memory retrieval, timeline retrieval, and no retrieval at all.
Lastly, although we include many recent frameworks as baselines, we failed to compare Theanine with MemoryBank (Zhong et al., 2024), a framework inspired by Ebbinghaus’s forgetting curve. This is because the time intervals between sessions in MSC and CC are either mostly measured in hours or not clearly specified (e.g., “a few months later”), whereas MemoryBank requires precise time intervals in days to apply the forgetting curve. Also, data used for MemoryBank focuses on Chinese clinical scenarios, making it not feasible for our study. However, we remain positive about applying such a mechanism to improve Theanine in our ongoing research.
10 Ethical Statements
LLMs might generate harmful, biased, offensive, sexual content. Authors avoid such content from appearing in this paper. We guarantee fair compensation for human evaluators from Amazon Mechanical Turk. We ensure an effective pay rate higher than 20\$ per hour based on the estimated time required to complete the tasks.
11 Acknowledgments
This work was partly supported by an IITP grant funded by the Korean Government (MSIT) (No. RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)). Jinyoung Yeo is the corresponding author (jinyeo@yonsei.ac.kr).
Appendix A Further Implementation Details
A.1 Cause-effect Commonsense Relations
We adopt and modify commonsense relations from Hwang et al. (2021) for our relation-aware memory linking. Below is the list of our commonsense relations $R$:
Changed: Events in A changed to events in B.
Cause: Events in A caused events in B.
Reason: Events in A are due to events in B.
HinderedBy: When events in B can be hindered by events in A, and vice versa.
React: When, as a result of events in A, the subject feels as mentioned in B.
Want: When, as a result of events in A, the subject wants events in B to happen.
SameTopic: When the specific topic addressed in A is also discussed in B.
Limited by the performance of retrievers, it is possible that an $m\in M_{a}$ does not have a relation, other than just textual overlap, with $m_{new}$. We address this by allowing the LLM to output None.
A.2 Algorithms for [Uncaptioned image]Theanine
The pseudo algorithms for Phase I and II are provided in Algorithm 1 and 2.
A.3 Computational Experiments
All computational experiments in this work are based on OpenAI API (OpenAI, 2024a). Thus, no computing infrastructure is required in this work.
Appendix B TeaFarm Evaluation
The overview of TeaFarm is illustrated in Figure 8.
Appendix C The [Uncaptioned image]TeaBag Dataset
We curate [Uncaptioned image]TeaBag, a dataset for TeaFarm evaluation on MSC and CC. TeaBag consists of:
  • 100 episodes of original conversations from Multi-Session Chat and Conversational Chronicles (session 1-5; 50 episodes from each dataset)
  • Two pairs of counterfactual QAs for each episode (200 pairs in total).
  • Two synthesized session 6 for each episode (thus 200 in total), each of which naturally leads to one of the counterfactual questions.
This dataset is made with GPT-4. The prompt for generation is in Appendix H. We expect future work to apply TeaBag to stress-test if their dialogue system can correctly reference past conversations.
TeaBag does not contain personally identifying information, as it is generated based on datasets where all contents are pure artificial creation, rather than contents collected from the real-world. Also, we have tried our best to confirm that this dataset does not contain any offensive content.
Appendix D G-Eval
G-Eval (Liu et al., 2023) is a framework using LLMs with chain-of-thoughts (CoT) and a form-filling paradigm, to assess the quality of models’ text generation. G-Eval with GPT-4 has been shown to generate evaluation results that highly align with human judgement (Liu et al., 2023; Kim et al., 2024b). Due to the high API cost for it, we conduct head-to-head comparisons between Theanine and 8 baselines using 100 episodes from MSC and CC (around 3,600 Theanine’s responses being compared in each comparison with one baseline), rather than the whole test sets. The prompts for G-Eval are in Figure 2021, and 22. We use SciPy to calculate p-values. [^10] https://scipy.org/
Appendix E Dataset-Specific G-Eval Results
We provide dataset-specific results of G-Eval in Table 4 and 5.
Appendix F Human Evaluation
We conduct human evaluation, with workers from Amazon Mechanical Turk (AMT). We construct the following three evaluations:
  • Appropriateness of relation-aware memory linking: In this evaluation, we ask the workers to judge whether they agree that the relation-aware linking is properly done for two given memories.
  • Helpfulness of context-aware timeline refinement: This evaluation requires the workers to determine if they agree that our context-aware refinement really tailors a raw timeline into a resource of useful information for generating the next response.
  • The quality of Theanine’s responses: Here, the workers are asked to judge if the responses of Theanine correctly refer to past conversations. After reading our responses and past memories, they choose whether the responses entail, contradict, or are neutral to past memories. To improve evaluation quality, we use GPT-4 to select responses for this specific evaluation based on past memories, addressing the fact that not every turn in the conversation requires previous information to generate the next response (In the other two evaluations, the samples are randomly selected).
Each data sample is judged by 3 different workers, and we report the results based on the majority rule. In the third evaluation, when every option (entailment, neutral, contradiction) gets one vote, we consider it neutral (13 samples in total). These human evaluations are conducted on 300 samples. The interfaces provided to AMT workers, which include detailed instructions for human evaluation, are shown in Figure 910, and 11.
Appendix G Empirical Examples
We provide several empirical examples of [Uncaptioned image]Theanine. Examples of relation-aware memory linking are in Figure 1213, and 14. Examples of utilizing refined timeline for response generation are in Figure 15.
Appendix H Prompts
The following are all prompts utilized in our study:
  • Relation-aware memory linking (Phase I-2): Figure 16.
  • Context-aware timeline refinement (Phase II-2): Figure 17.
  • Timeline-augmented Response generation (Phase III): Figure 18.
  • + Memory update (baseline): Figure 19.
  • RSum-LLM (baseline): We adopt the original prompt from Wang et al. (2023).
  • MemoChat (baseline): We adopt the original prompt from Lu et al. (2023).
  • COMEDY (baseline): We adopt the original prompt from Chen et al. (2024b).
  • G-Eval: Figure 20,  21, and  22.
  • Generating counterfactual QA in TeaFarm: Figure 23.
  • Generating session 6 in TeaFarm: Figure 24.
  • Evaluating model responses in TeaFarm: Figure 25.
Appendix I Terms for Use of Artifacts
We adopt the MSC and CC datasets from Xu et al. (2022a) and Jang et al. (2023), respectively. Both of these datasets are open-sourced for academic and non-commercial use. Our curated dataset, TeaBag, which will be released after acceptance, is open to academic and non-commercial use.
Algorithm 1 Memory Graph Construction (Phase I)
Require: Memory graph $G^{t}=(V^{t},E^{t})$
Require: New memories $M_{new}=\{m_{new1},...,m_{newN}\}$
Require: Set of relations $R=\{\texttt{Cause},\texttt{Reason},\texttt{Want},...,\texttt{SameTopic}\}$
Ensure: Memory graph $G^{t+1}=(V^{t+1},E^{t+1})$
1:$\Upsilon(m_{i},m_{j})=\begin{cases}r_{i,j},\text{if }m_{i}\text{ is assigned with }r_{i,j}\in R\text{ with }m_{j}\\ \texttt{None},\text{otherwise}\end{cases}$
2:$\Omega(V)=(\text{the most recent memory m}\in V)$
3:$E_{t+1}\leftarrow E_{t}$
4:for $m_{new}\in M_{new}$ do
5:     $M_{a}\leftarrow\{m_{i}\in V^{t}\mid m_{i}\text{ has top-{j} similarity with }m_{new}\}$
6:     $M_{a}^{*}\leftarrow\{m_{i}\in M_{a}\mid\Upsilon(m_{i},m_{new})=r\text{ for }r\in R\}$
7:     $\mathbb{C}\leftarrow\{C_{i}\mid C_{i}\text{ connected component of }G^{t}\text{ s.t. }\mathtt{V}(C_{i})\cap M_{a}^{*}\neq\emptyset\ \}$
8:     $M_{linked}\leftarrow\{\Omega(\mathtt{V}(C_{i})\cap M_{a}^{*})\mid C_{i}\in\mathbb{C}\}$
9:     $E_{new}\leftarrow\{\langle m_{i},\Upsilon(m_{i},m_{new}),m_{new}\rangle\mid m_{i}\in M_{linked}\}$
10:     $E_{t+1}\leftarrow E_{t+1}+E_{new}$
11:end for
12:$V^{t+1}\leftarrow V^{t}+M_{new}$
13:$G^{t+1}\leftarrow(V^{t+1},E^{t+1})$
14:return $G^{t+1}$
Algorithm 2 Timeline Retrieval and Timeline Refinement (Phase II)
Require: Memory graph $G=(V,E)$
Require: Dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$
Ensure: Collection of refined timelines $\mathbb{T}_{\Phi}$
1:$\Theta(V)=(\text{the oldest memory m}\in V)$
2:$M_{re}\leftarrow\{m_{i}\in V\mid m_{i}\text{ has top-{k} similarity with }D\}$
3:$\mathbb{C}_{re}\leftarrow\{C_{re}\mid C_{re}\text{ connected component of }G\text{ s.t. }\mathtt{V}(C_{re})\cap M_{re}\neq\emptyset\}$
4:$\mathbb{T}\leftarrow\{\}$
5:for $C_{re}\in\mathbb{C}_{re}$ do
6:     $m_{start}\leftarrow\Theta(\mathtt{V}(C_{re}))$
7:     $\mathcal{T}=\{\tau\subset C_{re}\mid\tau\text{ is a directed linear graph s.t.}m_{start},m_{re}\in\tau\ \land deg^{+}(\tau[-1])=0\}$
8:     $\mathbb{T}\leftarrow\mathbb{T}+\text{RandomSelection}(\mathcal{T})$
9:end for
10:$\mathbb{T}_{\Phi}\leftarrow\{\underset{\mathcal{T}_{\Phi}}{\text{argmax}}\,P_{\text{LLM}}(\mathcal{T}_{\Phi}|\mathcal{D},\tau)\mid\tau\in\mathbb{T}\}$
11:return $\mathbb{T}_{\Phi}$
Figure 8: The overview of TeaFarm Evaluation.
Datasets: Multi-session Chat (MSC)
Dialogue Session: Session 3 Session 4 Session 5 Avg.
[Uncaptioned image]Theanine Theanine’s Winning Percentage
          vs. Spe Mem Cst Spe Mem Cst Spe Mem Cst Spe Mem Cst
All Dialogue History 74% 79% 72% 78% 80% 74% 74% 76% 71% 75% 78% 72%
All Memories and $\mathcal{D}$ 71% 66% 64% 72% 69% 67% 70% 65% 63% 72% 66% 65%
     + Memory Update 59% 54%* 58% 58% 53% 58% 56%* 54%* 57%* 58% 54%* 58%
Memory Retrieval 68% 61% 65% 73% 67% 67% 72% 64% 65% 71% 64% 66%
     + Memory Update 56%* 53% 57% 61% 58% 61% 53%* 53% 53% 57% 55% 57%
RSum-LLM 44% 57% 68% 41% 54%* 67% 43% 53% 63% 42% 55% 66%
MemoChat 65% 75% 70% 64% 69% 72% 65% 74% 72% 65% 73% 71%
COMEDY 72% 64% 84% 72% 61% 84% 69% 60% 81% 71% 61% 83%
Table 4: Winning percentage of Theanine in head-to-head G-Eval (of all generated responses) with baselines. Tested with Multi-Session Chat (MSC). Color gray shows where ours loses more. Underline: p-value $<0.001$. *: $0.001<$ p-value $<0.05$.
Datasets: Conversation Chronicles (CC)
Dialogue Session: Session 3 Session 4 Session 5 Avg.
[Uncaptioned image]Theanine Theanine’s Winning Percentage
          vs. Spe Mem Cst Spe Mem Cst Spe Mem Cst Spe Mem Cst
All Dialogue History 69% 85% 83% 67% 82% 79% 71% 82% 82% 69% 83% 81%
All Memories and $\mathcal{D}$ 66% 62% 68% 66% 69% 70% 69% 71% 73% 67% 68% 70%
     + Memory Update 54%* 59% 62% 55%* 58% 59% 55%* 63% 64% 54% 60% 62%
Memory Retrieval 67% 64% 71% 70% 67% 68% 71% 66% 71% 69% 66% 70%
     + Memory Update 57% 59% 61% 55%* 57% 62% 57% 62% 66% 56% 60% 63%
RSum-LLM 31% 64% 69% 29% 60% 67% 32% 60% 68% 31% 61% 68%
MemoChat 56%* 73% 73% 57% 69% 71% 55%* 70% 70% 56% 70% 71%
COMEDY 59% 44% 73% 57% 46% 71% 55%* 46% 71% 57% 46% 72%
Table 5: Winning percentage of Theanine in head-to-head G-Eval (of all generated responses) with baselines. Tested with Conversation Chronicles (CC). Color gray shows where ours loses more. Underline: p-value $<0.001$. *: $0.001<$ p-value $<0.05$.
Figure 9: Interface for human evaluation regarding memory linking.
Figure 10: Interface for human evaluation regarding timeline refinement.
Figure 11: Interface for human evaluation regarding referencing past conversations in responses.
Figure 12: Examples of Relation-aware Memory Linking - 1.
Figure 13: Examples of Relation-aware Memory Linking - 2.
Figure 14: Examples of Relation-aware Memory Linking - 3.
Figure 15: Examples of Timeline Refinement and Response Generation.
Figure 16: The prompt for the Relation-aware memory linking.
Figure 17: The prompt for the context-aware timeline refinement.
Figure 18: The prompt for the timeline-augmented response generation.
Figure 19: The prompt for the memory updating mechanism in baselines (i.e., + Memory Update).
Figure 20: The prompt for the G-Eval: Specificity.
Figure 21: The prompt for the G-Eval: Memorability.
Figure 22: The prompt for the G-Eval: Consistency.
Figure 23: The prompt for generating counterfactual QA in TeaFarm.
Figure 24: The prompt for generating session 6 in TeaFarm.
Figure 25: The prompt for evaluating model response in TeaFarm.