### v2

# Towards Lifelong Dialogue Agents via Relation-aware Memory Construction and Timeline-augmented Response Generation

Abstract
To achieve lifelong human-agent interaction, dialogue agents need to constantly memorize perceived information and properly retrieve it for response generation (RG). While prior work focuses on getting rid of outdated memories to improve retrieval quality, we argue that such memories provide rich, important contextual cues for RG (e.g., changes in user behaviors) in long-term conversations. We present [Uncaptioned image]Theanine, a framework for LLM-based lifelong dialogue agents. Theanine discards memory removal and manages large-scale memories by linking them based on their temporal and cause-effect relation. Enabled by this linking structure, Theanine augments RG with memory timelines - series of memories representing the evolution or causality of relevant past events. Along with Theanine, we introduce TeaFarm, a counterfactual-driven evaluation scheme, addressing the limitation of G-Eval and human efforts in measuring memory-augmented dialogue agents. A supplementary video for Theanine and data for TeaFarm are at https://huggingface.co/spaces/ResearcherScholar/Theanine.
Towards Lifelong Dialogue Agents via Relation-aware Memory Construction and Timeline-augmented Response Generation
Kai Tzu-iunn Ong1thanks: KT Ong and N Kim are the co-first authors.    Namyoung Kim1    Minju Gwak1    Hyungjoo Chae1 Taeyoon Kwon1    Yohan Jo2    Seung-won Hwang2    Dongha Lee1    Jinyoung Yeo1 1Yonsei University, 2Seoul National University
1 Introduction
Autonomous agents based on large language models (LLMs) have made great progress in various domains, including response generation (Zhou et al., 2023; Kwon et al., 2024; Tseng et al., 2024), where agents ought to constantly keep track of both old and newly introduced information shared with users throughout their service lives (Irfan et al., 2024) and converse accordingly. To facilitate such lifelong interaction, studies have proposed to improve dialogue agents’ ability to memorize and correctly recall past information (e.g., discussed topics) in long-term, multi-session conversations.
Figure 1: Empirical examples of failed responses due to (a) absence of an important past event (“afraid of cruise ships”) on the timeline and (b) bias to the latest input. (c) is a response augmented with the memory timeline.
Figure 2: The overview of Refer to captionTheanine. Left: Linking new memories to the memory graph after finishing a dialogue session; Right: Memory timeline retrieval, refinement, and response generation in a new dialogue session.
A representative method is to compress past conversations into summarized memories and retrieve them to augment response generation (RG) in later encounters (Xu et al., 2022a; Lu et al., 2023). However, the growing span of memories can hinder retrieval quality as conversations accumulate. Although it, to some extent, can be solved by updating old memories (Bae et al., 2022; Zhong et al., 2024), this common practice may cause severe information loss. As shown in Figure 1 (a), an earlier memory on the timeline, an important persona (“afraid of ships”), is removed during memory update and causes improper RG. While using the large context windows of recent LLMs to process all dialogue history/memories is an option to prevent such information loss, [^1] For instance, GPT-4o and Llama 3.1 have context windows of 128K tokens (OpenAI, 2024a; MetaAI, 2024). this often leads to biased attention toward the latest user input (Figure 1 (b)), ignoring relevant contexts from the past (Liu et al., 2024). These findings highlight two main challenges towards lifelong dialogue agents - (i) Memory construction: how to effectively store large-scale past interactions without removing old memories? (ii) Response generation: within the growing memory span, how to identify relevant contextual cues for generating proper responses?
Motivated by these, we propose to address the above two challenges separately yet complementarily, by (i) discarding memory update to avoid information loss, and preserving relevant memories on the timeline in a linked structure; and (ii) retrieving the timeline as a whole to better catch relevant memories within the growing search span. We present [Uncaptioned image]Theanine, [^2] L-theanine is an amino acid found in green tea that has been linked to memory improvement (Nguyen et al., 2019). a framework for facilitating lifelong dialogue agents.
Starting from memory construction (Phase I), instead of stacking raw memory sentences as-is (Xu et al., 2022a), which may affect memory retrieval and also response quality due to the unstructured format of information (Mousavi et al., 2023; Chen et al., 2023), Theanine stores memories in a directed graph. In this graph, inspired by how humans naturally link new memories to existing ones of relevant events based on their relation (Bartlett, 1995), memories are linked using their temporal and cause-effect commonsense relations (Hwang et al., 2021). Supported by such linking structure, in memory retrieval for RG (Phase II-1), we go beyond conventional top-$k$ retrieval and further obtain the complete timelines to avoid missing out on important memories that have low textual overlap with current conversation (Tao et al., 2023). Lastly, to tackle the discrepancy between off-line memory construction and online deployment, Theanine uses an LLM to refine retrieved timelines (Phase II-2) based on current conversation (Madaan et al., 2024), such that they provide tailored information (Chae et al., 2023) for generating responses (Phase III). Our contributions are two-fold:
  • To achieve lifelong dialogue agents, we present [Uncaptioned image]Theanine, an LLM-based framework with a relation-aware memory graph and timeline-augmented response generation for long-term conversations. Theanine outperforms baselines across automatic, LLM-based, and human evaluations of RG. Also, we confirm that Theanine leads to better retrieval quality, and its intermediate procedures align with human preference. To our knowledge, we are the first to model the concept of timelines in memory management and RG.
  • The lack of golden mapping between conversations and reference memories poses a challenge in assessing memory-augmented agents. We present TeaFarm, a counterfactual-driven pipeline evaluating agent performance in referencing the past without human intervention.
2 Methodologies
We present [Uncaptioned image]Theanine, a framework for lifelong dialogue agents, inspired by how humans store and retrieve memories for conversations (Figure 2):
2.1 Memory Graph Construction (Phase I)
To manage large-scale memories and facilitate structured information for RG (Mousavi et al., 2023; Chen et al., 2023), we approach memory management using a memory graph $G$:
$G=(V,E)$ (1)
$V=\{m_{1},m_{2},...,m_{|V|}\}$ (2)
$m=(\textit{event},\textit{time})$ (3)
$E=\{\langle m_{i},r_{ij},m_{j}\rangle|m_{i},m_{j}\in V\land r_{ij}\in R\}$ (4)
$R=\{\texttt{Cause},\texttt{Reason},\texttt{Want},...,\texttt{SameTopic}\}$ (5)
In $G$, vertices $V$ are memories $m$ summarized from the conversations. Each memory $m=(\textit{event},\textit{time})$ consists of an event [^3] In this work, “event” denotes information perceived by the dialogue system, including things done/said by speakers and the acknowledgement of speaker personas. and the time it is formed (summarized). Each directed edge $e\in E$ between two connected $m$ indicates their temporal order and their cause-effect commonsense relation $r\in R$:
At the end of dialogue session $t$, [Uncaptioned image]Theanine starts linking each new memory $m_{new}$ summarized from session $t$ to the memory graph $G^{t}$.
Phase I-1: Identifying associative memories for memory linking.
Following how humans link new memories to existing ones that are related to a similar event/topic, i.e., the associative memories, Theanine starts by identifying these associative memories from the memory graph $G^{t}$.
Formally, given a newly-formed memory $m_{new}$ waiting to be stored, the associative memories $M_{a}$ of $m_{new}$ is defined as the set of $m_{i}\in G^{t}$ having top-$j$ text similarity with $m_{new}$ (i.e., $|M_{a}|=j$).
Phase I-2: Relation-aware memory linking.
Intuitively, we can link $m_{new}$ to $m\in M_{a}$ using edges that indicate their text similarity and chronological order, we find such simplified connection (e.g., “this happened $\rightarrow$ that similar event occurred”) can yield a context-poor graph that does not help response generation much (Section 4).
Humans, on the other hand, interpret events by considering the relation between them, such as “how does an event affect the other?” or “why did this person make that change?”. Therefore, we adopt a relation-aware memory linking, where an edge between two memories is encoded with their cause-effect commonsense relation $r\in R$, along w/ the temporal order. In practice, we adopt the commonly used relations defined by Hwang et al. (2021), including HinderedBy, Cause, Want, and 4 more (Appendix B.1).
We start by determining the relation between $m_{new}$ and each associative memory. Formally, for each pair of $m_{new}$ and $m\in M_{a}$, the LLM assigns a relation $r\in R$ based on their event, time and their origin conversations:
$M_{a}^{*}=\{m_{i}\in M_{a}\mid\Upsilon(m_{i},m_{new})\in R\}$ (6)
where $\Upsilon(\cdot,m_{new})\in R$ indicates that the given memory is assigned with an $r\in R$ with $m_{new}$, [^4] Limited by retrievers, an $m\in M_{a}$ may not have a relation with $m_{new}$. We thus allow the LLM to output “None”. and such assigned memories are defined as $M_{a}^{*}$.
We then proceed to link $m_{new}$ to the graph. We first locate every connected component $C_{i}\subset G^{t}$ that contains at least one $m\in M_{a}^{*}$, as shown in Figure 3 (a) and (b):
=$\mathbb{C}=\{C_{i}\subset G^{t}\mid\mathtt{V}(C_{i})\cap M_{a}^{*}\neq\emptyset\ \}$ (7)
where $\mathbb{C}$ is the collection of those $C$ and $\mathtt{V}(\cdot)$ represents “vertices in”. Then, we link $m_{new}$ to the most recent [^5] Simply linking $m_{new}$ to all $m\in M_{a}^{*}$ costing 25% more API cost for linking without leading to better response. $m\in M_{a}^{*}$ in each $C_{i}\subset\mathbb{C}$ (Figure 3 (c)). Memories $M_{linked}$ that are linked to $m_{new}$ is defined as follows:
$M_{linked}=\{\Omega(\mathtt{V}(C_{i})\cap M_{a}^{*})\mid C_{i}\subset\mathbb{C}\}$ (8)
where $\Omega(\cdot)$ indicates “the most recent memory in”.
Figure 3: Locating memories to be linked to $m_{new}$.
Linking all memories from session $t$ to $G^{t}$, we then obtain a new memory graph $G^{t+1}$. The pseudo algorithm for Phase I is in Algorithm 1.
2.2 Timeline Retrieval and Timeline Refinement (Phase II)
Thanks to the constructed memory graph, Theanine can proceed to augment RG with timelines of relevant events, addressing the information loss in conventional memory management (Figure 1). With $G^{t+1}$, Theanine performs the following steps for RG in session $t+1$:
Preparation: Top-k memory retrieval.
During the conversation, using the current dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ of $n$ utterances $u$ as query, we retrieve top-$k$ memories $M_{re}=\{m_{re1},...,m_{rek}\}$.
Phase II-1: Retrieving and untangling raw memory timelines.
We wish to also access memories centered around $M_{re}$. Formally, given $m_{re}\in M_{re}$, we further collect the connected component $C_{re}\subset G^{t+1}$ that contains $m_{re}$ via the linked structure.
Since this collection of memories (i.e., $C_{re}$) can be “tangled up” together (i.e., connected in a complex manner) due to the graph structure, we proceed to untangle it into several memory timelines, each representing a series of events about $m_{re}$ that starts out similarly yet branches into slightly different development. For that, we first locate the earliest memory in $C_{re}$ as a starting point $m_{start}$ for all timelines, as shown in Figure 4 (left).
$m_{start}=\Theta(\mathtt{V}(C_{re}))$ (9)
where $\Theta$ indicates “the oldest memory in”.
Figure 4: Extracting raw memory timelines $\tau$ from the retrieved connected component $C_{re}$.
Next, starting from $m_{start}$, we untangle memories by tracing through future direction and extract every possible linear graph containing $m_{re}$ (two in Figure 4) from $C_{re}$, until reaching an endpoint $\tau[-1]$ with an out-degree of 0 (i.e., $deg^{+}(\tau[-1])=0$), which means no directed edge goes out from it). Each of them is considered a raw memory timeline $\tau$, demonstrating a version of the evolution of $m_{re}$ and its relevant events:
$\begin{split}\mathcal{T}={}&\{\tau\subset C_{re}\mid\tau\smash{{\text{ is a directed linear}}}\\ &\smash{{\text{ graph s.t. }}}m_{start},m_{re}\in\tau\ \\ &\land deg^{+}(\tau[-1])=0\}\end{split}$ (10)
We then sample $n$ raw timelines $\tau$ from $\mathcal{T}$ [^6] Based on our experiments, we empirically set $n$ to 1. Repeating777“Repeating” is used to explain the algorithm from the perspective of one $m_{re}$. In practice, $M_{re}$ are processed together, although processing them 1-by-1 yields the same result. Phase II-1 for all retrieved top-$k$ memories, we collect a set of retrieved raw memory timelines $\mathbb{T}=\cup\,\mathcal{T}$, where $|\mathbb{T}|=k^{*}n$.
Phase II-2: Context-aware timeline refinement.
Although we have constructed the memory graph using temporal and commonsense relations to improve its informativeness, directly applying retrieved timelines for RG can be suboptimal (RQ3, Section 4), because graph construction does not take current conversation into consideration, i.e., they are constructed off-line.
In this phase, Theanine tackles such a discrepancy between off-line memory construction and online deployment (i.e., ongoing conversation) via a context-aware timeline refinement. Motivated by how LLMs can refine their previous generation (Madaan et al., 2024). We leverage LLMs to refine raw timelines into a rich resource of information crafted for the current conversation, by removing redundant information or highlighting information that can come in handy. Formally, given the current dialogue $\mathcal{D}$ and retrieved raw timelines $\mathbb{T}$, an LLM tailors $\tau\in\mathbb{T}$ into refined timelines $\mathbb{T}_{\Phi}$:
$\mathbb{T}_{\Phi}=\{\underset{\tau_{\Phi}}{\text{argmax}}\,P_{\text{LLM}}(\tau_{\Phi}|\mathcal{D},\tau)\mid\tau\in\mathbb{T}\}$ (11)
All refined timelines $\mathbb{T}_{\Phi}$ are then used to augment the response generation. We provide the pseudo algorithm for Phase II in Algorithm 2.
2.3 Timeline-augmented Response Generation (Phase III)
Now, Theanine utilizes the refined timelines for RG. Formally, given $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$ and $\mathbb{T}_{\Phi}$, an LLM generates a next response $\bar{u}_{t+1}$:
$\bar{u}_{n+1}=\underset{u_{n+1}}{\text{argmax}}\,P_{\text{LLM}}({u}_{n+1}|\mathcal{D},\mathbb{T}_{\Phi})$ (12)
3 Experimental Setups
Datasets: Multi-session Chat (MSC) Conversation Chronicles (CC)
Methods / Metrics Bleu-4 Rouge-L Mauve BertScore Bleu-4 Rouge-L Mauve BertScore
All Dialogue History 1.65 14.89 9.06 86.28 4.90 21.56 26.47 88.13
All Memories & Current Context $\mathcal{D}$ 1.56 14.89 10.62 86.23 4.41 20.06 38.16 88.02
+ Memory Update (Bae et al., 2022) 1.55 14.77 9.28 86.20 4.34 20.34 34.84 88.03
Memory Retrieval (Xu et al., 2022a) 1.92 15.49 11.16 86.47 4.93 20.63 33.06 88.07
+ Memory Update (Bae et al., 2022) 1.67 15.30 13.71 86.39 4.46 20.19 34.28 88.02
Rsum-LLM (Wang et al., 2023) 0.75 11.53 2.45 84.91 0.98 11.42 2.28 85.59
MemoChat (Lu et al., 2023) 1.42 13.51 7.72 85.96 2.31 15.87 15.12 87.08
COMEDY (Chen et al., 2024b) 1.06 12.79 7.27 85.29 1.70 13.57 1.95 85.90
[Uncaptioned image]Theanine (Ours) 1.80 15.37 18.62 86.70 6.85 22.68 64.41 88.58
Table 1: Automatic evaluation of response quality (average of sessions). Session-specific results are in Appendix F.
3.1 Datasets of Long-term Conversations
There are few datasets of long-term, multi-session conversations. First, Multi-Session Chat (MSC) (Xu et al., 2022a), which is built upon Persona-Chat (Zhang et al., 2018) by extending its conversations to multiple (five) sessions. Soon after MSC, DuLeMon (Xu et al., 2022b) and CareCall (Bae et al., 2022) are proposed for long-term conversations in Mandarin and Korean. Recently, Jang et al. (2023) release a new dataset, Conversation Chronicles (CC). Different from MSC, speakers in CC are augmented with relationships, such as “employee and boss”. Apart from the above open-domain datasets, there is also Psychological QA, [^8] https://www.xinli001.com/ which addresses long-term conversations in Mandarin regarding clinical scenarios.
We use MSC and CC for evaluation. We focus on English conversations, so we do not adopt DuleMon, CareCall, and Psychological QA and leave multilingual and clinical settings to future work.
3.2 Baselines
To evaluate Theanine, besides naive baselines using all past dialogues or memories, we include:
Memory Retrieval. Following Xu et al. (2022a), we use a retriever to retrieve memories relevant to the current dialogue context to augment RG.
Memory Update. We use LLMs to run a widely used updating algorithm proposed by Bae et al. (2022) at the end of each session. The algorithm includes functionalities such as Change, Replace, Delete, Append, and more (see Appendix H).
RSum-LLM. An LLM-only generative method, which recursively summarizes and updates the memory pool and accordingly generates responses (w/o a retrieval module) (Wang et al., 2023).
MemoChat. Proposed by Lu et al. (2023), it uses LLMs’ CoT reasoning ability to (i) conclude important memories from past conversations in a structured topic-summary-dialogue manner, (ii) select memories, and (ii) generate responses.
COMEDY. Proposed by Chen et al. (2024b), it uses LLMs to summarize session-level memories, compresses all of them into short events, user portraits (behavioral patterns, emotion, etc.) and user-bot relation, and selects them to augment RG.
3.3 Models and Implementation Details
Large language models. In all experiments including baselines, we adopt gpt-3.5-turbo-0125 (OpenAI, 2023) for (i) memory summarization (Table 4), (ii) memory update, and (iii) response generation. We set the temperature to $0.75$.
Retrievers. We use text-embedding-3-small (OpenAI, 2024b) to calculate text similarity. In the identification of top-$j$ associative memories (Phase I-1) and top-$k$ memory retrieval (Phase II), we set $j$ and $k$ to 3. For “Memory Retrieval” baseline, we retrieve 6 memories following Xu et al. (2022a).
Dialogue sessions. We use sessions 3-5 of MSC and CC for evaluations, as all methods are almost the same in sessions 1-2 (no memory to update).
4 Evaluation Scheme 1: Automatic and Human Evaluations
To evaluate [Uncaptioned image]Theanine’s responses in long-term conversations, we follow common practices and conduct 3 types of evaluations: (i) Automatic evaluations; (ii) G-Eval (Liu et al., 2023), an LLM-based framework commonly used to evaluate LMs’ generation; (iii) human evaluation. We now present the main findings (details, prompts, and interfaces of evaluations in Scheme 1 are in Appendix E):
(Finding 1) Theanine outperforms baselines in response generation.
Table 1 reports agent performance in RG regarding both overlap-based and embedding-based metrics: Bleu-4 (Papineni et al., 2002), Rouge-L (Lin, 2004), Mauve (Pillutla et al., 2021), and BertScore (Zhang et al., 2020). Theanine, in both datasets, achieves superior response quality than diverse baselines. Despite slightly lower scores compared to Memory Retrieval in overlap-based metrics (i.e., B-4 and R-L) in MSC, Theanine largely outperforms it in terms of embedding-based metrics. Interestingly, including ours, methods w/o memory update generally yield higher scores, justifying our proposal towards an update-free memory management for lifelong dialogue agents.
Settings / Metrics B-4 R-L Mauve Bert
[Uncaptioned image]Theanine (Ours) 4.32 19.03 41.52 87.64
w/o Relation-aware Linking 4.07 18.58 39.69 87.57
w/o Timeline Refinement 4.03 18.82 41.34 87.66
Broken Down, Shuffled Timeline 4.15 18.70 38.49 87.61
Memory Retrieval 3.43 18.06 22.11 87.27
Table 2: Performance of our ablations (avg. of datasets).
(Finding 2 & 3) All phases contribute to performance; retrieving the timeline as a whole brings large improvement over conventional retrieval.
To gain deeper insights into our design, we show Theanine’s performance when relation-awareness in linking (Phase I-2) and Timeline Refinement (Phase II-2) are ablated. Also, to objectively assess whether Theanine’s retrieval (i.e., retrieving the timeline as a whole) improves retrieval quality, we include a setting where retrieved timelines are broken down into randomly ordered events such that retrieved memories during RG are in the same format as conventional top-$k$ retrieval.
In Table 2, we observe a ranking regarding contribution to performance: relation-aware linking $>$ retrieving timeline as a whole $>$ timeline refinement. We thus confirm the efficacy of constructing a memory graph with causal relations. Moreover, utilizing such a structure to collect timelines of relevant events yields better RG than conventional retrieval, despite the smaller $k$ (3 vs. 6) in initial retrieval. Refining timelines shows smaller performance gains, suggesting room for improvement in applying them for RG. We leave it to future work.
(Finding 4) Humans and G-Eval both reveal that Theanine retrieves memories that are more helpful for response generation.
Beyond agent responses, we further investigate how different memory construction methods affect retrieved memories’ helpfulness in augmenting RG. Given the same current dialogues as queries for retrieval, Figure 5 (left; G-Eval; right: Human Eval) shows head-to-head comparisons (ours vs baselines) regarding whose retrieved memories benefit response generation more. We observe higher win rates of Theanine in all comparisons, especially in human evaluations, suggesting that ours can facilitate a more helpful memory-augmentation for RG.
Figure 5: Human- (right) and machine-based (left) head-to-head comparisons between ours and baselines regarding the helpfulness of retrieved memories.
(Finding 5) Humans confirm that Theanine yields responses better entailing past interactions.
Now that the helpfulness of Theanine’s retrieved memories is validated, we proceed to investigate whether such helpful memories can help us facilitate reliable lifelong dialogue agents.
For that, we further ask a group of workers to specifically judge whether agent responses entail, contradict, or are neutral to the past via majority voting. In Figure 6, we find Theanine not only leading to a small number of contradictory responses (4%) but also showing the largest percentage (68%; out of 100) of responses that entail past conversations, significantly outperforming baselines. We argue that it is because our timeline-based approach elicits memories better at representing past interactions between speakers, thus leading to responses that more directly align with the past, which is important for dialogue agents to maintain long-term intimacy with users (Adiwardana et al., 2020). Also, such entail, non-contradictory nature of Theanine’s responses highlights its potential for applications in specialized domains, such as personalized agents for clinical scenarios, where entailment between agent responses and users’ past information (e.g., electrical health records or previous consulting sessions) is crucial for diagnostic decison-making (Tseng et al., 2024).
Figure 6: Human evaluations regarding to what extent the agent responses entail past conversations.
As a side note, Memory Update yields fewer contradictory responses (2%), indicating a potential trade-off between (i) removing outdated memories to prevent contradiction and (ii) preserving them to get richer information for RG (Kim et al., 2024a).
(Finding 6) Humans agree with Theanine’s intermediate procedures.
As reported in Figure 7, judges largely agree (92%) that Theanine properly assigns cause-effect relations to linked memories, explaining its contribution to performance. Also, they agree that timeline refinement successfully elicits more helpful information (100%; 100 samples in total) for RG. Examples of Theanine’s phases and RG are in Appendix G.
Figure 7: Human evaluation of our intermediate phases.
5 Evaluation Scheme 2: TeaFarm – a Counterfactual-driven Evaluation Pipeline for Long-term Conversations
Evaluating memory-augmented agents in long-term conversations is non-trivial due to the unavailability of ground-truth mapping between current conversations and correct memories for retrieval. Although we may resort to G-Eval by feeding evaluator LLMs (e.g., GPT-4) the entire past history and prompt it to determine whether a response correctly recalls the past, the evaluation can be largely limited by the performance of the evaluator LLM itself (Kim et al., 2024b).
To overcome this, along with Theanine, we present TeaFarm, a human-free counterfactual-driven pipeline for evaluating memory-augmented response generation in long-term conversations.
5.1 Testing Dialogue Agents’ Memory via Counterfactual Questions
In TeaFarm, we proceed to “trick” dialogue agents into generating incorrect responses, and agents must correctly reference past conversations to avoid being misled by us. Specifically, we talk to the dialogue agent while acting as if a non-factual statement is true (thus counterfactual). Figure 8 presents some examples of counterfactual questions and the corresponding facts.
Figure 8: Examples of counterfactual questions.
In practice (Figure 10), when we want to evaluate an agent that has been interacting with the user for sessions, we first (1) collect all past conversations and summarize them session by session. Then, we (2) feed a question generator LLM [^9] We apply GPT-4 (gpt-4) with a temperature of $0.75$. the collected summaries in chronological order such that it can capture the current stage of each discussed event, e.g., “Speaker B does not own a car”, and (3) generate counterfactual questions from the perspective of both speakers (and the correct answers). After that, we (4) kick off (i.e., simulate) a new dialogue session, chat for a while, then (5) naturally ask the counterfactual question, and (6) assess the correctness of its response. The overview figure, prompts, and synthesized data for TeaFarm are in Appendix CH, and D, respectively.
5.2 TeaFarm Results
Settings / Datasets MSC CC Avg.
Memory Retrieval 0.16 0.19 0.18
+ Memory Update 0.16 0.19 0.18
RSum-LLM 0.04 0.08 0.06
MemoChat 0.09 0.15 0.12
COMEDY 0.06 0.18 0.12
[Uncaptioned image]Theanine 0.18 0.24 0.21
      w/o Relation-aware Linking 0.17 0.20 0.19
      w/o Timeline Refinement 0.16 0.19 0.18
Table 3: Success rates (SRs) of correctly recalling the past and not being fooled by the counterfactual questions in TeaFarm (tested with 200 questions).
In Table 3, Theanine shows higher SR than baselines, especially in CC. Ablations perform slightly worse than the original, again proving the efficacy of relation-aware linking and timeline refinement. Surprisingly, all settings have low SRs, qualifying TeaFarm as a proper pipeline for stress-testing dialogue agents in long-term conversations.
Interestingly, baselines using retrievers (same as Theanine) show superior performance than settings only relying on LLMs (i.e., RSum-LLM, MemoChat, and COMEDY). This, unexpectedly, supports our efforts in developing a new paradigm of memory management in the era of LLMs. [^10] Memory update does not affect Memory Retrieval’s performance. We believe it is because counterfactual questions are made to counter the newest stage of each event. The removal of older memories thus does not have much impact.
6 Further Analyses and Discussions
Cost efficiency.
A concern of Theanine is the API cost. Regardless, we argue that it is competitive when both performance and cost are taken into account. Figure 9 plots response quality (Mauve score) against the API cost. [^11] Calculated Based on session 5, which involves most memories for management. We use Mauve for its stronger correlation with humans over diverse metrics (Pillutla et al., 2021). We find Theanine and all ablations not only outperform all baselines but also lie on the Pareto frontier, indicating an efficient cost-performance trade-off. This suggests Theanine’s value when performance is prioritized over API costs.
Figure 9: Cost-performance comparisons.
Growing span of memories.
Another inquiry left might be whether the growing span of memory will eventually hinder memory retrieval in Theanine if there ever are hundreds of sessions. Although this can be a serious issue for conventional memory retrieval, we presume that it will be mitigated to some extent in Theanine because: (i) We retrieve relevant memories as a whole in the form of timelines. This serves as a safety net in scenarios where an important memory is missed out in top-$k$ retrieval – it might still be retrieved along with the timeline; (ii) We perform timeline refinements to further adjust retrieved timelines based on current dialogue context such that it provided tailored information for the response generator. This acts as a second insurance against sub-optimal retrieval.
7 Related Work
Long-term Conversations.
Since MSC, there are several studies on long-term conversations: Bae et al. (2022) train a classifier to update old memories in phone call scenarios. As we enter the era of LLMs, Wang et al. (2023) and Li et al. (2024) leverage LLMs to write and update memories for RG. Apart from LLMs’ power, human behaviors also foster methods in this field. For example, Zhong et al. (2024) apply humans’ forgetting curve to make memories that have been discussed stay in the memory pool longer. Recently, Maharana et al. (2024) also adopt the concept of timelines. However, in their work, a timeline is a fixed, pre-defined series of events (potentially unrelated), which simply serve as a user profile for synthesizing dialogue data. By contrast, in our work, a timeline is built with relevant events, which are dynamically linked and retrieved as the conversation goes on, serving as useful cues for response generation.
Memory-augmentation for personalized dialogue agents.
The trend of long-term interaction with autonomous agents promotes their adaptation for personalized needs (Chen et al., 2024a, c). As a pioneer, Xu et al. (2022b) train a persona extractor to create user-based memories. However, training personalized agents for long-term use can be non-trivial due to the lack of data (Tseng et al., 2024). As a solution, Kim et al. (2024a) apply commonsense models and LLMs to augment existing long-term data with high-quality persona sentences; Chen et al. (2024b) present a training-free LLM-based framework that extracts user behaviors from conversations for personalized RG. Upon the success of LLMs, Theanine leverages them to build memory timelines. These timelines represent the development of interactions and lead to responses that better entail speaker information, establishing Theanine’s potential for personalized agents.
8 Conclusions
This paper presents a first-ever timeline-based framework for memory-augmented dialogue agents in long-term conversations. Comprehensive evaluations show that [Uncaptioned image]Theanine facilitates more beneficial augmentation, leading to responses that are closer to ground-truths and more aligned with speakers’ past interactions. Its effectiveness is further confirmed in TeaFarm, a QA-based pipeline we propose to address the limitation of G-Eval and human efforts in assessing memory-augmented RG. We expect our novel approach to serve as a new foundation towards lifelong dialogue agents.
9 Limitations
First, the amount of dialogue sessions in this study is limited to five due to the lack of longer open-domain English datasets. As we mentioned in Section 6, we presume that Theanine’s effectiveness can still hold true to some degree in longer conversations. Yet, we do acknowledge the need to apply additional modules that directly address the growing span of dialogue history/memories, such as introducing the summarize-then-compress paradigm in COMEDY (Chen et al., 2024b) to compress session-level summaries into a combined short user/event description.
Second, although we include many recent frameworks as baselines, we failed to compare Theanine with MemoryBank (Zhong et al., 2024), a framework inspired by Ebbinghaus’s forgetting curve. This is because the time intervals between sessions in MSC and CC are either mostly measured in hours or not clearly specified (e.g., “a few months later”), whereas MemoryBank requires precise time intervals in days to apply the forgetting curve. Also, data used for MemoryBank focuses on Chinese clinical scenarios, making it not feasible for our study. However, we remain positive about applying such a mechanism to improve Theanine in our ongoing research.
Lastly, API-based LLMs may introduce risks such as privacy issues. A possible solution is to apply Theanine to small open-source LMs for secure, local usage. While there exist challenges in data collection, one may achieve this by (i) collecting synthesized conversations with GPT-generated user profiles, (ii) running Theanine on these data, and (iii) using the outputs of each phase to train student LMs (i.e., distillation from teacher LLMs).
10 Ethical Statements
LLMs might generate harmful, biased, offensive, sexual content. Authors avoid such content from appearing in this paper. We guarantee fair compensation for human evaluators from Amazon Mechanical Turk. We ensure an effective pay rate higher than 20\$ per hour based on the estimated time required to complete the tasks.
11 Acknowledgments
This work was partly supported by an IITP grant funded by the Korean Government (MSIT) (No. RS-2020-II201361, Artificial Intelligence Graduate School Program (Yonsei University)). Jinyoung Yeo is the corresponding author (jinyeo@yonsei.ac.kr).
Appendix A Appendix Contents
  • Appendix B.1: Cause-effect Commonsense Relations.
  • Appendix B.2: Algorithms for [Uncaptioned image]Theanine.
  • Appendix B.3: Implementation Details on Computational Experiments.
  • Appendix C: TeaFarm Evaluation.
  • Appendix D: The [Uncaptioned image]TeaBag Dataset.
  • Appendix E Details on Evaluation Scheme 1 (G-Eval and Human Evaluations).
  • Appendix F: Session-specific Evaluation Results.
  • Appendix G: Empirical Examples.
  • Appendix H: Prompts Used in This Work.
  • Appendix I: Further Analyses and Discussions.
  • Appendix J: Terms for Use of Artifacts.
Appendix B Further Implementation Details
B.1 Cause-effect Commonsense Relations
We adopt and modify commonsense relations from Hwang et al. (2021) for our relation-aware memory linking. Below is the list of our commonsense relations $R$:
Changed: Events in A changed to events in B.
Cause: Events in A caused events in B.
Reason: Events in A are due to events in B.
HinderedBy: When events in B can be hindered by events in A, and vice versa.
React: When, as a result of events in A, the subject feels as mentioned in B.
Want: When, as a result of events in A, the subject wants events in B to happen.
SameTopic: When the specific topic addressed in A is also discussed in B.
Limited by the performance of retrievers, it is possible that an $m\in M_{a}$ does not have a relation, other than just textual overlap, with $m_{new}$. We address this by allowing the LLM to output None.
B.2 Algorithms for [Uncaptioned image]Theanine
The pseudo algorithms for Phase I and II are provided in Algorithm 1 and 2.
B.3 Implementation Details on Computational Experiments
All computational experiments in this work are based on OpenAI API (OpenAI, 2024a). Thus, no computing infrastructure is required in this work.
Appendix C TeaFarm Evaluation
The overview of TeaFarm is illustrated in Figure 10.
Appendix D The [Uncaptioned image]TeaBag Dataset
As a byproduct of TeaFarm, we curate [Uncaptioned image]TeaBag, a dataset for TeaFarm evaluation on MSC and CC. TeaBag consists of:
  • 100 episodes of original conversations from Multi-Session Chat and Conversational Chronicles (session 1-5; 50 episodes from each dataset)
  • Two pairs of counterfactual QAs for each episode (200 pairs in total).
  • Two synthesized follow-up conversations (i.e., session 6) for each episode (thus 200 in total), each of which naturally guides the conversation from session 5 towards one of the counterfactual questions.
This dataset is made with GPT-4. The prompt for generation is in Appendix H. We expect future work to apply TeaBag to stress-test if their dialogue system can correctly reference past conversations.
TeaBag does not contain personally identifying information, as it is generated based on datasets where all contents are pure artificial creation, rather than contents collected from the real-world. Also, we have tried our best to confirm that this dataset does not contain any offensive content.
For the overview of data collection, please refer to step 1-4 of TeaFarm (Figure 10).
Appendix E Details on Evaluation Scheme 1
We perform evaluations using sessions 3-5 from MSC and CC, as all settings are almost identical before the end of session 2, due to the fact that there is no memory to update before then.
The test sets of MSC and CC contain over 500 and 20,000 episodes of conversations, where each episode has 5 dialogue sessions, yielding 1.2M turns of responses in total. Due to the limited budget for generation (both baselines and ours), when not specified, we sample 50 episodes from each dataset for experiments in this paper (around 3.6K conversational turns in total).
E.1 G-Eval
G-Eval (Liu et al., 2023) is a framework using LLMs with chain-of-thoughts (CoT) and a form-filling paradigm to assess the quality of models’ text generation. G-Eval with GPT-4 has been shown to generate evaluation results that highly align with human judgement (Liu et al., 2023; Kim et al., 2024b) and thus has been widely applied in many LM-based projects. We conduct G-Eval on 5 episodes.
The prompt for evaluating the helpfulness of retrieved memories is in Figure 23. We use SciPy to calculate p-values. [^12] https://scipy.org/
E.2 Human Evaluation
We conduct human evaluation, with workers from Amazon Mechanical Turk (AMT). We construct the following three evaluations:
  • Appropriateness of relation-aware memory linking: In this evaluation, we ask the workers to judge whether they agree that the relation-aware linking is properly done for two given memories. The interface provided to AMT workers, which includes detailed instructions for human evaluation, is shown in Figure 11.
  • Helpfulness of context-aware timeline refinement: This evaluation requires the workers to determine if they agree that our context-aware refinement really tailors a raw timeline into a resource of useful information for generating the next response. The interface provided to AMT workers, which includes detailed instructions for human evaluation, is shown in Figure 12.
  • The quality of responses: Here, the workers are asked to judge if the responses correctly refer to past conversations. After reading our responses and past memories, they choose whether the responses entail, contradict, or are neutral to past memories. To improve evaluation quality, we use GPT-4 to select responses for this specific evaluation based on past memories, addressing the fact that not every turn in the conversation requires previous information to generate the next response (In the other two evaluations, the samples are randomly selected). The interface provided to AMT workers, which includes detailed instructions for human evaluation is shown in Figure 13.
  • The helpfulness of retrieved memories: Given the same dialogue context, human workers are asked to select a memory that is more helpful for generating a next response from ours’ and a baseline’s retrieval. The interface provided to AMT workers, which includes detailed instructions for human evaluation is shown in Figure 14
Each data sample is judged by 3 different workers, and we report the results based on the majority rule. In the third evaluation, when every option (entailment, neutral, contradiction) gets one vote, we consider it neutral (13 samples in total). These human evaluations are conducted on 100 conversational turns.
Appendix F Session-specific Evaluation Results
We provide session-specific results for automatic evaluations in Table 5.
Appendix G Empirical Examples
We provide several empirical examples of [Uncaptioned image]Theanine. Examples of relation-aware memory linking are in Figure 1516, and 17. Examples of utilizing refined timeline for response generation are in Figure 18.
Appendix H Prompts
The following are all prompts utilized in our study:
  • Relation-aware memory linking (Phase I-2): Figure 19.
  • Context-aware timeline refinement (Phase II-2): Figure 20.
  • Timeline-augmented Response generation (Phase III): Figure 21.
  • Memory Update (baseline): Figure 22.
  • RSum-LLM (baseline): We adopt the original prompt from Wang et al. (2023).
  • MemoChat (baseline): We adopt the original prompt from Lu et al. (2023).
  • COMEDY (baseline): We adopt the original prompt from Chen et al. (2024b).
  • G-Eval: The prompt for evaluating the helpfulness of retrieved memories is in Figure 23.
  • Generating counterfactual QA in TeaFarm: Figure 24.
  • Generating session 6 in TeaFarm: Figure 25.
  • Evaluating model responses in TeaFarm: Figure 26.
Appendix I Further Analyses and Discussions
Memory Summarization.
At the end of each session, we use ChatGPT (gpt-3.5-turbo-0125) to summarize the conversation into memory sentences. We conduct examinations on such conversation-memory summarization using 100 randomly sampled sessions from MSC and CC to make sure the quality of raw memories is acceptable. The result is in Table 4.
Memories that … No Can’t judge Yes
Contain faulty statements 90% 9% 1%
Miss important statements 95% 4% 1%
Table 4: Human evaluation of conversation-to-memory summarization in Theanine.
Appendix J Terms for Use of Artifacts
We adopt the MSC and CC datasets from Xu et al. (2022a) and Jang et al. (2023), respectively. Both of these datasets are open-sourced for academic and non-commercial use. Our curated dataset, TeaBag, which will be released after acceptance, is open to academic and non-commercial use.
Algorithm 1 Memory Graph Construction (Phase I)
Require: Memory graph $G^{t}=(V^{t},E^{t})$
Require: New memories $M_{new}=\{m_{new1},...,m_{newN}\}$
Require: Set of relations $R=\{\texttt{Cause},\texttt{Reason},\texttt{Want},...,\texttt{SameTopic}\}$
Ensure: Memory graph $G^{t+1}=(V^{t+1},E^{t+1})$
1:$\Upsilon(m_{i},m_{j})=\begin{cases}r_{i,j},\text{if }m_{i}\text{ is assigned with }r_{i,j}\in R\text{ with }m_{j}\\ \texttt{None},\text{otherwise}\end{cases}$
2:$\Omega(V)=(\text{the most recent memory m}\in V)$
3:$E_{t+1}\leftarrow E_{t}$
4:for $m_{new}\in M_{new}$ do
5:     $M_{a}\leftarrow\{m_{i}\in V^{t}\mid m_{i}\text{ has top-{j} similarity with }m_{new}\}$
6:     $M_{a}^{*}\leftarrow\{m_{i}\in M_{a}\mid\Upsilon(m_{i},m_{new})=r\text{ for }r\in R\}$
7:     $\mathbb{C}\leftarrow\{C_{i}\mid C_{i}\text{ connected component of }G^{t}\text{ s.t. }\mathtt{V}(C_{i})\cap M_{a}^{*}\neq\emptyset\ \}$
8:     $M_{linked}\leftarrow\{\Omega(\mathtt{V}(C_{i})\cap M_{a}^{*})\mid C_{i}\in\mathbb{C}\}$
9:     $E_{new}\leftarrow\{\langle m_{i},\Upsilon(m_{i},m_{new}),m_{new}\rangle\mid m_{i}\in M_{linked}\}$
10:     $E_{t+1}\leftarrow E_{t+1}+E_{new}$
11:end for
12:$V^{t+1}\leftarrow V^{t}+M_{new}$
13:$G^{t+1}\leftarrow(V^{t+1},E^{t+1})$
14:return $G^{t+1}$
Algorithm 2 Timeline Retrieval and Timeline Refinement (Phase II)
Require: Memory graph $G=(V,E)$
Require: Dialogue context $\mathcal{D}=\{u_{i}\}_{i=1}^{n}$
Ensure: Collection of refined timelines $\mathbb{T}_{\Phi}$
1:$\Theta(V)=(\text{the oldest memory m}\in V)$
2:$M_{re}\leftarrow\{m_{i}\in V\mid m_{i}\text{ has top-{k} similarity with }D\}$
3:$\mathbb{C}_{re}\leftarrow\{C_{re}\mid C_{re}\text{ connected component of }G\text{ s.t. }\mathtt{V}(C_{re})\cap M_{re}\neq\emptyset\}$
4:$\mathbb{T}\leftarrow\{\}$
5:for $C_{re}\in\mathbb{C}_{re}$ do
6:     $m_{start}\leftarrow\Theta(\mathtt{V}(C_{re}))$
7:     $\mathcal{T}=\{\tau\subset C_{re}\mid\tau\text{ is a directed linear graph s.t.}m_{start},m_{re}\in\tau\ \land deg^{+}(\tau[-1])=0\}$
8:     $\mathbb{T}\leftarrow\mathbb{T}+\text{RandomSelection}(\mathcal{T})$
9:end for
10:$\mathbb{T}_{\Phi}\leftarrow\{\underset{\mathcal{T}_{\Phi}}{\text{argmax}}\,P_{\text{LLM}}(\mathcal{T}_{\Phi}|\mathcal{D},\tau)\mid\tau\in\mathbb{T}\}$
11:return $\mathbb{T}_{\Phi}$
Figure 10: The overview of TeaFarm Evaluation.
Datasets: Multi-session Chat (MSC) & Conversation Chronicals (CC)
Session: Session 3 Session 4 Session 5
Methods / Metrics B-4 R-L Mauve Bert B-4 R-L Mauve Bert B-4 R-L Mauve Bert
All Dialogue History 3.13 18.04 17.34 87.17 3.17 17.96 18.54 87.12 3.53 18.69 17.42 87.31
All Memories & Current Context $\mathcal{D}$ 2.69 17.29 28.30 87.10 3.10 17.38 22.52 87.06 3.16 17.75 22.35 87.21
+ Memory Update (Bae et al., 2022) 2.80 17.51 22.92 87.11 2.88 17.24 21.22 86.99 3.16 17.90 22.04 87.24
Memory Retrieval (Xu et al., 2022a) 3.44 18.33 24.68 87.30 3.38 17.55 21.95 87.17 3.46 18.31 19.70 87.33
+ Memory Update (Bae et al., 2022) 3.10 18.08 25.02 87.24 2.99 17.37 25.97 87.10 3.11 17.78 20.99 87.28
Rsum-LLM (Wang et al., 2023) 0.83 11.30 2.45 85.25 0.87 11.35 2.32 82.20 0.90 11.78 2.33 85.30
MemoChat (Lu et al., 2023) 1.88 14.83 14.56 86.57 1.81 14.27 10.57 86.43 1.91 14.96 9.13 86.56
COMEDY (Chen et al., 2024b) 1.14 12.80 4.74 85.53 1.57 13.18 5.16 85.56 1.42 13.56 3.94 85.69
[Uncaptioned image]Theanine (Ours) 4.21 19.21 45.53 87.70 4.42 18.63 37.84 87.52 4.34 19.23 41.18 87.70
Table 5: Session-specific results of agent performance in response generation.
Figure 11: Interface for human evaluation regarding memory linking.
Figure 12: Interface for human evaluation regarding timeline refinement.
Figure 13: Interface for human evaluation regarding referencing past conversations in responses.
Figure 14: Interface for human evaluation regarding the helpfulness of retrieved memories.
Figure 15: Examples of Relation-aware Memory Linking - 1.
Figure 16: Examples of Relation-aware Memory Linking - 2.
Figure 17: Examples of Relation-aware Memory Linking - 3.
Figure 18: Examples of Timeline Refinement and Response Generation.
Figure 19: The prompt for the Relation-aware memory linking.
Figure 20: The prompt for the context-aware timeline refinement.
Figure 21: The prompt for the timeline-augmented response generation.
Figure 22: The prompt for the memory updating mechanism in baselines (i.e., + Memory Update).
Figure 23: The prompt for the G-Eval: Helpfulness of Retrieved Memories.
Figure 24: The prompt for generating counterfactual QA in TeaFarm.
Figure 25: The prompt for generating session 6 in TeaFarm.
Figure 26: The prompt for evaluating model response in TeaFarm.